# Can LLMs Understand Time Series Anomalies?

### v2

@@ 4.2 Prompting Strategies
We incorporate two main prompting techniques in our investigation: Zero-Shot and Few-Shot Learning (FSL) and Chain of Thought (CoT) For FSL, we examine the LLM’s ability to detect anomalies without any examples (zero-shot) and with a small number of labeled examples (few-shot). For CoT, we implement example in-context CoT templates, guiding the LLM through a step-by-step reasoning process. Our template prompts the LLM to: (1) Recognize and describe the general time series pattern (e.g., periodic waves, increasing trend) (2) Identify deviations from this pattern (3) Determine if these deviations constitute anomalies based on the dataset’s normal behaviors.
@@ 4.2.1 Input Representation
We examine several text encoding strategies to enhance the LLM’s comprehension of time series data:(1) Original: Raw time series values presented as rounded space-separated numbers. (2) CSV: Time series data formatted as CSV (index and value per line, comma-separated), inspired by Jin et al. (2024b). (3) Prompt as Prefix (PAP): Including key statistics of the time series (mean, median, trend) along with the raw data, as suggested by Jin et al. (2024a). (4) Token per Digit (TPD): Splitting floating-point numbers into space-separated digits (e.g., 0.2462 4 6) to circumvent the OpenAI tokenizer’s default behavior of treating multiple digits as a single token, following Gruver et al. (2023). This strategy only improve the performance of models that apply byte-pair encoding (BPE) tokenization, see Appendix C Observation 7.
@@
We utilize Matplotlib to generate visual representations of the time series data. These visualizations are then provided to multimodal LLMs capable of processing image inputs. Since LLMs have demonstrated strong performance on chart understanding tasks (Shi et al., 2024), they are expected to identify anomaly regions’ boundaries from the visualized time axis.
@@ 5.1 Experiment Setup
Table 1: Variants and their corresponding namecodes
@@
Figure 2: Reflexive/Reflective, Top 3 Affi-F1 prompt variant per mode (Ignoring (dys)calc variants), See Table 1 for variant name codes and Appendix A.4 for details
@@
The LLMs generate anomaly periods that can be converted to binary labels and do not output anomaly scores. Therefore, we report precision, recall, and F1-score metrics. However, these scores treat time series as a cluster of points without temporal order and can give counterintuitive results, as illustrated in Figure 3. To address this, we also report affinity precision and affinity recall as defined in Huet et al. (2022). We calculate the affinity F1 score as the harmonic mean of affi-precision and affi-recall, and it is the main metric we use to evaluate the hypotheses.
@@
As our goal is not to propose a new anomaly detection method but rather to test hypotheses about LLMs for better understanding, we use simple baselines for sanity check, see Appendix C Observation 8. We use Isolation Forest (Liu et al., 2008) and Thresholding.
@@ 5.2 Experiment Results
In this section, we discuss the hypotheses that align with our observations and those we can confidently reject. Detailed numbers can be found in Appendix D. The 0-to-1 y-axis of the figures represents the affinity F1 score, where higher values indicate better performance. We defer some non-essential observations to Appendix C to focus on the hypotheses.
Retained Hypothesis 1 on CoT Reasoning No evidence is found that explicit reasoning prompts (via Chain of Thought) improved LLMs’ performance in time series analysis.
@@
Figure 4: Clean/Noisy, Top 3 Affi-F1 variants per noise level
@@
Figure 5: Calc/DysCalc, Top 3 Affi-F1 variants per mode
Figure 6: Vision/Text, Top 3 Affi-F1 variants per modality
Figure 7: Subsampled/Original, 0-shot raw text vs 30% text
Rejected Hypothesis 3 on Arithmetic Reasoning The LLM’s understanding of time series is not related to its ability to perform arithmetic calculations.
@@
Figure 8: Flat Trend/Trend, Top 3 Affi-F1 variants per dataset
@@ 6 Conclusion
In this paper, we conducted a comprehensive investigation into Large Language Models’ (LLMs) understanding of time series data and their anomaly detection capabilities. Our findings challenge several assumptions prevalent in current literature, highlighting the need for rigorous empirical validation of hypotheses about LLM behavior. Our key findings include LLM’s visual advantage, limited reasoning, non-human-like processing, and model-specific capabilities. These insights have important implications for the design of future LLMs and the development of anomaly detection systems. For instance, our results suggest that LLMs do not effectively detect visual frequency anomalies, so vision-LLM-based anomaly detection systems shall leverage Fourier analysis before feeding the data to the LLM To improve performance. Our model-specific findings suggest that model selection and possible ensemble methods are crucial for designing LLM-based anomaly detection systems.
@@ Appendix C Other Findings
Observation 7 on BPE tokenization Only OpenAI GPT with the BPE tokenization can occasionally benefits from Token-per-Digit representation of input.
@@
Observation 8 on LLM Performance LLMs are reasonable choices for zero-shot time series anomaly detection, giving superior performance compared to traditional methods in many cases.
@@
Observation 9 on Optimal Text Representation Across all text representation methods, no single method consistently outperforms the others.

### v3

@@ 4.2 Prompting Strategies
We incorporate two main prompting techniques in our investigation: Zero-Shot and Few-Shot Learning (FSL) and Chain-of-Thought (CoT). For FSL, we examine the LLM’s ability to detect anomalies without any examples (zero-shot) and with a small number of labeled examples (few-shot). For CoT, we implement example in-context CoT templates, guiding the LLM through a step-by-step reasoning process. Our template prompts the LLM to: (1) Recognize and describe the general time series pattern (e.g., periodic waves, increasing trend) (2) Identify deviations from this pattern (3) Determine if these deviations constitute anomalies based on the dataset’s normal behaviors.
@@ 4.2.1 Input Representation
We examine several text encoding strategies to enhance the LLM’s comprehension of time series data:(1) Original: Raw time series values presented as rounded space-separated numbers. (2) CSV: Time series data formatted as CSV (index and value per line, comma-separated), inspired by Jin et al. (2024). (3) Prompt as Prefix (PAP): Including key statistics of the time series (mean, median, trend) along with the raw data, as suggested by Jin et al. (2023). (4) Token per Digit (TPD): Splitting floating-point numbers into space-separated digits (e.g., 0.2462 4 6) to circumvent the OpenAI tokenizer’s default behavior of treating multiple digits as a single token, following Gruver et al. (2023). This strategy only improve the performance of models that apply byte-pair encoding (BPE) tokenization, see Appendix C Observation 8.
@@
We utilize Matplotlib to generate visual representations of the time series data. These visualizations are then provided to multimodal LLMs capable of processing image inputs. Since LLMs have demonstrated strong performance on chart understanding tasks (Shi et al., 2024), they are expected to identify anomaly regions’ boundaries from the visualized time axis.=
@@ 5.1 Experiment Setup
Table 1: Variants and their corresponding namecodes, see Appendix A.4 for details
@@
We choose synthetic data instead of real-world data due to: (1) known label quality issues in public benchmarks (Wu & Keogh, 2021), (2) need for precise anomaly type isolation, (3) requirement for textual anomaly descriptions to enable CoT prompting. Still, our experiments on real-world Yahoo S5 dataset Laptev & Amizadeh (2015) (see Appendix D) show consistent findings.
@@
The LLMs generate anomalous intervals that can be converted to binary labels and do not output anomaly scores. Therefore, we report precision, recall, and F1-score metrics. However, these scores treat time series as a cluster of points without temporal order and can give counterintuitive results, as illustrated in Figure 2. To address this, we also report affinity precision and affinity recall as defined in Huet et al. (2022). We calculate the affinity F1 score as the harmonic mean of affi-precision and affi-recall, and it is the main metric we use to evaluate the hypotheses. We rely on variants of F1 because the LLM yields discrete intervals, not an anomaly score. Metrics like Volume Under the Surface (Paparrizos et al., 2022) require a ranking or continuous score.
@@
Figure 3: Reflexive (prompt that induces reasoning) / Reflective (prompt asks for direct answer), Top 3 Affi-F1 prompt variant per mode, See Table 1 for variant name codes.
@@
As our goal is not to propose a new anomaly detection method but rather to test hypotheses for better understanding, we use simple baselines for sanity check, see Appendix C Observation 9 for direct performance comparison with baselines. We use Isolation Forest (Liu et al., 2008) and Thresholding.
@@ 5.2 Experiment Results
In this section, we discuss the hypotheses that align with our observations and those we can confidently reject. Detailed numbers can be found in Appendix D. The 0-to-1 y-axis of the figures represents the affinity F1 score, where higher values indicate better performance. We focus on the 6 key hypotheses and defer other findings to Appendix C.
Retained Hypothesis 1 on CoT Reasoning No evidence is found that explicit reasoning prompts via CoT improve LLMs’ performance in time series analysis.
@@
Figure 4: Clean (original time series) / Noisy (time series with minimal injected noise), Top 3 Affi-F1 variants per noise level
@@
Figure 5: Calc (prompt with correct arithmetic example) / DysCalc (incorrect example), Top 3 Affi-F1 variants per mode
Figure 6: Vision (prompt with visualized time series) / Text (raw numerical prompt), Top 3 Affi-F1 variants per modality
Figure 7: Subsampled (time series subsampled to be shorter) / Original, 0-shot raw text vs 30% text
Rejected Hypothesis 3 on Arithmetic Reasoning The LLMs’ understanding of time series is not related to its ability to perform arithmetic calculations.
@@
Figure 8: Flat Trend (see above for an example) / Trend (trend may reverse during anomalies), Top 3 Affi-F1 variants per dataset
@@
Model Variations.
Individual models exhibit distinctly different behaviors in zero‐shot and few‐shot anomaly detection. For instance, GPT’s performance with longer sequences do not degrade as much as other models, and models like Qwen handle visual inputs far more successfully. Detailed comparisons and analyses of these model‐specific effects are provided in Appendix C. These observations underscore that LLMs’ performance in time series tasks can depend heavily on factors like training data, parameter count, and fine‐tuning strategies.
@@ 6 Conclusion
In this paper, we conducted a comprehensive investigation into Large Language Models’ (LLMs) understanding of time series data and their anomaly detection capabilities. Our findings challenge several assumptions prevalent in current literature, highlighting the need for rigorous empirical validation of hypotheses about LLM behavior. Our key findings include LLMs’ visual advantage, limited reasoning, non-human-like processing, and model-specific capabilities. These insights have important implications for the design of future LLMs and the development of anomaly detection systems. For instance, our results suggest that LLMs do not effectively detect visual frequency anomalies, so vision-LLM-based anomaly detection systems shall leverage Fourier analysis before feeding the data to the LLM to improve performance. Our model-specific findings suggest that model selection and possible ensemble methods are crucial for designing LLM-based anomaly detection systems.
@@
Acknowledgement
This work was supported in part by the U.S. Army Research Office under Army-ECASE award W911NF-07-R-0003-03, the U.S. Department Of Energy, Office of Science, IARPA HAYSTAC Program, and NSF Grants #2205093, #2146343, #2134274, CDC-RFA-FT-23-0069, DARPA AIE FoundSci and DARPA YFA. We thank Dr. Eamonn Keogh for the helpful suggestions.
@@ 1. Point Anomalies
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 9: Example time series from the Point Anomalies dataset, with anomalies regions highlighted in blue.
@@ 2. Range Anomalies
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 10: Example time series from the Range Anomalies dataset, with anomalies regions highlighted in blue.
@@ 3. Trend Anomalies
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 11: Example time series from the Trend Anomalies dataset, with anomalies regions highlighted in blue.
@@ 4. Frequency Anomalies
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 12: Example time series from the Frequency Anomalies dataset, with anomalies regions highlighted in blue.
@@ 5. Noisy Point Anomalies
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 13: Example time series from the Noisy Point Anomalies dataset, with anomalies regions highlighted in blue.
@@ 6. Noisy Trend Anomalies
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 14: Example time series from the Noisy Trend Anomalies dataset, with anomalies regions highlighted in blue.
@@ 7. Noisy Frequency Anomalies
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 15: Example time series from the Noisy Frequency Anomalies dataset, with anomalies regions highlighted in blue.
@@ 8. Flat Trend Anomalies
We conducted a human validation study with 5 participants detecting 20 subtle anomalies. Human detection rate was 0/20 vs. 17.2/20 for obvious anomalies, confirming the perceptual challenge.
@@
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 16: Example time series from the Flat Trend Anomalies dataset, with anomalies regions highlighted in blue.
9. Yahoo S5
The Webscope S5 dataset Laptev & Amizadeh (2015) is a widely-used and publicly accessible benchmark for anomaly detection. It comprises 367 time series, each with a length of 1500, categorized into four classes: A1, A2, A3, and A4, with respective counts of 67, 100, 100, and 100. Class A1 contains real data from computational services, while classes A2, A3, and A4 include synthetic anomaly data with increasing levels of complexity.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 17: Example time series from the Yahoo S5 dataset, with anomalies(mostly individual points) highlighted in red.
@@ Appendix C Other Findings
Hypothesis 7 on Model‐Family Variations LLMs’ time series understanding are consistent across different model families.
This hypothesis stems from the tendency in recent literature (Tang et al., 2024; Tan et al., 2024; Zeng et al., 2023) to generalize findings about time series understanding across all LLMs based on experiments with a limited set of models, typically GPT and LLaMA variants. This approach implies that different model families’ comprehension of time series data is universally consistent, varying primarily with the number of parameters. However, unlike in NLP tasks where specific models excel in areas like translation or mathematics (Cobbe et al., 2021), there’s a lack of understanding regarding specialized skills in time series analysis across different LLM models. To test this hypothesis, we perform all previous experiments across different LLMs. If the hypothesis holds, we should see previous conclusions either consistently validated or invalidated across all models.
Observation 8 on BPE tokenization Only OpenAI GPT with the BPE tokenization can occasionally benefits from Token-per-Digit representation of input.
@@
Observation 9 on LLM Performance LLMs are reasonable choices for zero-shot time series anomaly detection, giving superior performance compared to traditional methods in many cases.
@@
Observation 10 on Optimal Text Representation Across all text representation methods, no single method consistently outperforms the others.
@@ Appendix D Full Experiment Results
Table 10: Yahoo S5 dataset
PRE REC F1 affi affi affi
PRE REC F1
Gemini-1.5-Flash 0shot-Text 0.18 3.16 0.26 16.10 18.50 16.46
0shot-Text-S0.3 20.21 22.73 20.10 31.04 30.41 29.91
0shot-Text-S0.3-CSV 3.82 3.95 3.82 13.80 10.71 11.60
0shot-Text-S0.3-PAP 16.48 16.67 16.51 26.52 23.85 24.53
0shot-Text-S0.3-TPD 14.97 16.62 15.11 28.32 27.09 27.00
0shot-Vision 30.44 45.17 31.74 54.82 57.25 55.77
1shot-Text-S0.3 20.41 21.93 20.40 31.39 30.66 30.32
1shot-Vision 28.54 31.06 28.56 45.27 43.10 43.64
GPT-4o-Mini 0shot-Text 2.74 4.04 2.50 15.26 14.68 14.40
0shot-Text-S0.3 36.02 37.42 36.06 41.38 41.57 41.26
0shot-Text-S0.3-CSV 18.28 18.48 18.29 27.40 25.31 25.86
0shot-Text-S0.3-PAP 25.62 25.70 25.63 29.31 28.48 28.70
0shot-Text-S0.3-TPD 4.17 4.65 4.22 11.98 11.00 11.20
0shot-Vision 50.88 57.84 51.31 71.42 74.07 72.19
1shot-Text-S0.3 39.74 41.03 39.73 46.59 45.90 45.76
1shot-Vision 17.22 18.07 17.18 32.81 34.18 32.61
Internvlm-76B 0shot-Text 30.42 37.91 30.54 38.31 42.25 39.53
0shot-Text-S0.3 46.08 47.28 46.10 48.13 48.47 48.13
0shot-Text-S0.3-CSV 6.30 8.49 6.34 16.59 14.92 15.32
0shot-Text-S0.3-PAP 53.72 54.52 53.76 55.43 55.64 55.49
0shot-Text-S0.3-TPD 14.80 17.51 14.86 19.96 21.27 20.29
0shot-Vision 12.42 29.87 13.52 35.19 40.91 37.41
1shot-Text-S0.3 50.17 52.32 50.21 52.12 53.23 52.46
1shot-Vision 1.84 11.19 2.47 20.62 23.15 20.96
Isolation-Forest 0shot 2.07 29.26 3.25 16.75 33.18 21.94
Qwen 0shot-Text 0.00 0.00 0.00 0.00 0.00 0.00
0shot-Text-S0.3 0.00 0.00 0.00 0.00 0.00 0.00
0shot-Text-S0.3-CSV 0.00 0.00 0.00 0.00 0.00 0.00
0shot-Text-S0.3-PAP 19.89 19.89 19.89 19.91 19.90 19.91
0shot-Text-S0.3-TPD 2.18 2.18 2.18 2.19 2.21 2.20
0shot-Vision 3.10 19.21 3.40 16.19 25.37 19.53
1shot-Text-S0.3 0.00 0.00 0.00 0.00 0.00 0.00
1shot-Vision 3.39 5.68 3.24 18.37 21.61 19.22
Threshold 0shot 3.99 21.99 5.51 16.40 31.80 21.03
Appendix E Mathematical Formulations of Hypotheses 2 and 3
E.1 Hypothesis 2: Repetition Bias and Periodicity Detection
Let $f(t)$ be a time series and $T(f)$ be its tokenized representation in an LLM’s vocabulary space $\mathcal{V}$. We define:
1) Perfect periodicity: $f(t+P)=f(t)$ for some period $P>0$
2) Noisy periodicity: $f(t+P)=f(t)+\epsilon(t)$ where $\epsilon(t)\sim\mathcal{N}(0,\sigma^{2})$ and $\sigma\ll\min_{t,t^{\prime}}|f(t)-f(t^{\prime})|$
Note that while noisy periodicity is defined on numerical values, the tokenization process $T(\cdot)$ maps these values to discrete tokens, making perfect periodicity in token space impossible for noisy periodic signals.
Let $A(f)$ be the LLM’s anomaly detection accuracy on time series $f$ and $\mathcal{D}_{P}$ be the set of all periodic time series with period $P$. Given there exists an optimal anomaly detector $B$, whose accuracy is $B^{*}(f)$, that can achieve near-perfect accuracy on both perfect and noisy periodic signals. The hypothesis states:
For any $f_{1},f_{2}\in\mathcal{D}_{P}$, if $T(f_{1})$ exhibits token-level periodicity and $T(f_{2})$ does not, then:
$E[A(f_{1})]\gg E[A(f_{2})]$
while the optimal detector maintains consistent performance:
$B^{*}(f_{1})\approx B^{*}(f_{2})\approx 1$
This formulation suggests that LLMs’ performance difference is due to token-level repetition bias rather than the inherent complexity of the anomaly detection task, as a properly designed detector can achieve near-perfect performance on both cases.
E.2 Hypothesis 3: Arithmetic Ability and Pattern Recognition
Let $M$ be an LLM and $M^{\prime}$ be the same LLM with impaired arithmetic ability.
1) Define arithmetic ability $\alpha(M)$ as accuracy on basic arithmetic tasks:
$\alpha(M)=E_{x,y}[\mathbb{1}[M(\text{"What is }x+y\text{?"})=x+y]]$
2) Define reasoning ability $\rho(M)$ as accuracy on non-arithmetic reasoning tasks:
$\rho(M)=E_{q\in\mathcal{Q}}[\mathbb{1}[M(q)=\text{correct}]]$
where $\mathcal{Q}$ is a set of logical reasoning questions.
3) Obtain $M^{\prime}$ by training $M$ on incorrect arithmetic examples while preserving reasoning:
$\begin{cases}\alpha(M^{\prime})\ll\alpha(M)\\ \rho(M^{\prime})\approx\rho(M)\end{cases}$
4) Hypothesis holds if:
$E_{f\in\mathcal{D}}[A_{M}(f)]\gg E_{f\in\mathcal{D}}[A_{M^{\prime}}(f)]$
where $\delta$ is small and dataset $\mathcal{D}$ is arbitrary. To falsify the hypothesis, we show the difference is negligble or reversed for certain datasets.